OpenStudy (zenmo):

Find the solution of the differential equation that satisfies the given initial condition. (Check my work please :D ).

OpenStudy (zenmo):

\[y'tanx = 9a+y, y(\pi/3)=9a, o < x < \pi/2\]a is a constant.

OpenStudy (zenmo):

Multiply both sides by dx/ (tanx)(a+y) \[\frac{ dy }{ a+y }=\frac{ 9dx }{ tanx }\]\[\int\limits_{}^{}\frac{ 1 }{ a+y }dy=9\int\limits_{}^{}cotxdx\]\[\ln \left| a+y \right|=9\ln \left| sinx \right|+K\] Replace K with lnC with positive constant \[\ln \left| a+y \right|=9\ln(sinx)+lnC\]\[\ln \left| a+y \right|=9\ln(Csinx)\]

OpenStudy (zenmo):

Raise power of e to both sides, \[\left| a+y \right|=9Csinx\] When x=pi/3, y = 9a \[\left| a+9a \right|= 9C \frac{ \sqrt{3} }{ 2 }\]\[\frac{ 2 }{ \sqrt{3} }(\frac{ 1 }{ 9})(10\left| a \right|)= C\]\[\frac{ 20 }{ 9\sqrt{3} }\left| a \right|=C\]\[\left| a+y \right|=9(\frac{ 20 }{ 9\sqrt{3} }\left| a \right|)sinx\]\[y = \frac{ 20 }{ \sqrt{3} }asinx-a\]

OpenStudy (zenmo):

Got this answer wrong, what do?

ganeshie8 (ganeshie8):

9a+y is not same as 9(a+y)

OpenStudy (mww):

yup just a simple factor problem.

OpenStudy (zenmo):

\[\ln \left| a+y \right|=9\ln \left| sinx \right|+C\]\[\left| a+y \right|=e^{9\ln \left| sinx \right|+C} = e^{9\ln \left| sinx \right|}*e^C\]\[\left| a+y \right|=9\left| sinx \right|e^C, K = \pm e^C\]\[a+y = 9sinx K => y = 9sinxK-a\]\[x = \pi/3, y = 9a\]\[a + 9a = K \sin \frac{ \pi }{ 3 }, 10a = \frac{\sqrt{3}}{ 2 }K, K = \frac{ 20a }{ \sqrt{3} }\] Thus, \[y = \frac{ 180a }{ \sqrt{3} }sinx-a\] Answer is still wrong, at what part did I mess up on with 9a+y not being the same as 9(a+y)?

OpenStudy (mww):

@Zenmo The very first part, sadly: y′tanx=9a+y dytanx = (9a+y)dx dy/(9a+y) = cot(x) dx In(9a+y) = In|sin(x)| + K = ln(sin(x)) + K for some constant C since sin is positive over (0,pi/2) ***Note carefully we never factored the 9 out to start with! The nine belongs to the constant a, not the variable y so there is no need to factor is out!) 9a+y = exp(ln(sinx)+K) 9a + y = sin(x).exp(K) y = exp(K). sin(x) - 9a = Csin(x) - 9a for some positive constant C Sub in x = pi/3, y = 9a 9a = Csin(pi/3) - 9a C = 18a/(sqrt(3)/2) = 36a/sqrt(3) So y = 36a/(sqrt(3) * sin(x) - 9a = 9a(4sin(x)/(sqrt(3)) - 1)